# Spontaneous phase locking in a broad-area semiconductor laser

### v1

@@ D.2 Simulations
In this section we present simulations of a heterodyning signal in order to demonstrate how the phase fluctuations can be extracted from it and to understand the effect of applying a frequency filter. We numerically calculate a heterodyning beat note $I_{m}(t)$ as in Eq. (21) with a time step of $\tau=20$ ps where we choose constant mode amplitudes $A_{m}=A_{r}=1$, a beat note frequency $\Delta\omega_{mr}=2$ GHz, and linewidths of $\Gamma_{m}=100$ MHz and $\Gamma_{r}=1$ MHz for the lasing mode and the reference laser, respectively. The phases $\phi_{m}$ and $\phi_{r}$ of lasing mode and reference laser are calculated as in Eq. (7). The beat note $I_{m}$ is a sinusoidal signal with phase modulated by the phase fluctuations of the two lasing modes.
@@
Figure 15: Analysis of simulated beat note signal. a, Amplitude of the analytic signal of the simulated beat note without filter ($A_{mr}$, blue) and with filter ($A_{mr}^{(f)}$, red). b, Phase extracted from the simulated beat note without filter ($\phi_{mr}$, blue) and with filter ($\phi_{mr}^{(f)}$, red) as well as the original phase ($\phi_{m}-\phi_{r}$, green). c, Original phase fluctuations ($\delta_{\tau}(\phi_{m}-\phi_{r})$, green) and phase fluctuations extracted with filter ($\delta_{\tau}\phi_{mr}^{(f)}$, red).
@@
First, we apply the Hilbert transform directly to $I_{m}(t)$ and extract the amplitude $A_{mr}(t)$ and the phase $\phi_{mr}(t)$ of the beat note as in Eqs. (23)–(25). The amplitude shown as blue line in Fig. 15 exhibits small fluctuations around $2$ even though it is expected to be constant and equal to $2$. Figure 15b compares the phase $\phi_{mr}(t)$ (blue) extracted from the analytic signal to the original phase $\phi_{m}-\phi_{r}$ (green) used to calculate the beat note, which agree very well. This demonstrates that the procedure outlined above allows us to determine the phase fluctuations of the lasing modes from the beat notes, even though there are small deviations between extracted amplitude and phase compared to the simulated ones.
@@
However, in practice the measured heterodyne signals contain several beat notes as well as contributions from the BAL itself, hence a relatively narrow filter must be applied around the individual beat notes in order to separate them from the other signals. To test the effect of filtering, we apply a $300$ MHz wide filter [cf. Eq. (33) below] to the simulated beat note $I_{m}(t)$. Then we calculate the analytic signal of the filtered time trace $I_{m}^{(f)}$ and extract amplitude $A_{mr}^{(f)}$ and phase $\phi_{mr}^{(f)}$ as above, which are shown in red in Fig. 15. The amplitude of the filtered signal shows severe fluctuations, including dips going to almost zero. These fluctuations are artifacts induced by the filter: in the event of very fast phase changes, for example at $46$ ns in Fig. 15b, the instantaneous frequency of the beat note can leave the frequency range of the filter, thus reducing the amplitude of the filtered signal. This unfortunately means that our heterodyne analysis does not allow us to determine a possible fast amplitude dynamics of the lasing modes since any genuine fluctuations of the amplitude will be obscured by the spurious amplitude fluctuations due to the filter.
@@
The phase (Fig. 15b) is also affected by the filter. First, it is smoothed because very fast phase fluctuations are filtered out. Hence the phase fluctuations of the filtered signal $\delta_{\tau}\phi_{mr}^{(f)}$ (red line in Fig. 15c) are much smaller than the original ones (green line) and are also smooth, not randomly fluctuating like the original ones. Second, strong dips in the amplitude lead to spurious fluctuations of the phase (phase slips) like at $46$ ns where $\phi_{mr}^{(f)}$ no longer follows the original phase. These phase slips manifest in strong extrema of the phase fluctuations. Once the amplitude has recovered (from $55$ ns on), the filtered phase once again follows the original one, but with an offset. In order to remove these spurious phase fluctuations induced by the filter, $4$ ns long windows around local minima of the amplitude are cut out from the phase fluctuation time traces before calculating their correlation.
@@
Finally we investigate how the data analysis procedure affects the correlation measurement. To this end, we calculate the beat note $I_{n}$ of another lasing mode $n$ with the reference laser. Lasing modes $n$ and $m$ have the same linewidth of $100$ MHz, but we now introduce a correlation $\rho_{0}$ of their phase fluctuations. The phase fluctuations of mode $n$ are calculated from the random variable $z_{n}$ [see Eq. (7)] given by
@@
where $z_{m}$ and $z_{3}$ are independent Gaussian random variables with unit variance. By construction, $\rho_{z_{m},z_{n}}=\rho_{0}$. We calculate the Pearson correlation coefficient $\rho_{m,n}^{(f)}$ of the filtered phase fluctuations $\delta_{\tau}\phi_{mr}^{(f)}$ and $\delta_{\tau}\phi_{nr}^{(f)}$, after additionally removing the parts near local minima of the amplitude as described above. The resulting correlations $\rho_{m,n}^{(f)}$ are compared to the actual correlations $\rho_{0}$ of the simulated phase fluctuations in Fig. 16. We find that for small $\rho_{0}$, the extracted PFCs are $\rho_{m,n}^{(f)}\simeq\rho_{0}/2$ and remain significantly lower even for high $\rho_{0}$. In conclusion, the analysis procedure used to measure the PFCs yields values that are systematically lower than the actual correlations, which partially explains the relatively low PFCs observed experimentally. Nonetheless, our simulations demonstrate that we obtain qualitatively correct information about possible phase correlations between different lasing modes.
@@ D.3 Analysis of experimental data
where $\Gamma_{m}$ is the full width at half maximum (FWHM) of the beat note averaged over $x$, $\Gamma_{m}=\langle\Gamma_{m,x}\rangle_{x}$. Here, the FWHM $\Gamma_{m,x}$ at position $x$ is obtained from the smoothed RF-spectrum of the heterodyne signal,
$P_{hd}(f,x)=\mathcal{M}\{|\tilde{U}(f,x)|^{2}\}$ (35)
where $\mathcal{M}$ is a moving average with $10$ MHz width, by determining the $-3$ dB points around the peak at $f_{m}$. Positions for which the FWHM could not be properly determined due to, e.g., low signal strength, are ignored when calculating $\Gamma_{m}$. A total width of $3\Gamma_{m}$ for the filter is used as a compromise between the requirements of having a large bandwidth and removing noise and interfering signals as much as possible. For a Lorentzian line shape, the filter thus contains $79.5\%$ of the total RF-power of the beat note signal. In practice, the filter width is of the order of a few $100$ MHz.
@@
where $\mathrm{sgn}$ is the sign function. In the following we consider the phase fluctuations
@@
where $\tau=20$ ps is the sampling step of the oscilloscope. The phase fluctuations $\delta_{\tau}\phi_{mr}^{(f)}(t,x)$ are the difference between the phase fluctuations of the lasing mode and that of the reference laser, where the latter are negligible compared to the former since the linewidth of the reference laser is very small compared to that of the lasing modes in our case.
@@
It should be noted that the standard deviation of $\delta_{\tau}\phi_{mr}^{(f)}(t,x)$ is significantly reduced compared to that of the actual phase fluctuations due to the filter that is applied. Furthermore, very fast phase fluctuations lead to a high instantaneous frequency that causes the signal to (partially) leave the frequency domain of the filter that is applied (see Fig. 15). In such cases, $\delta_{\tau}\phi_{mr}^{(f)}$ shows spurious values, and the instantaneous amplitude $A_{mr}^{(f)}$ exhibits a strong dip, as discussed in the previous section. To remove these filter artifacts, $4$ ns long pieces of $\delta_{\tau}\phi_{mr}^{(f)}$ are cut around pronounced dips in the amplitude. Finally, the Pearson correlation coefficient $\rho_{m,n}^{(f)}(x)$ between these reduced versions of $\delta_{\tau}\phi_{m}^{(f)}(t,x)$ and $\delta_{\tau}\phi_{n}^{(f)}(t,x)$ is calculated. As a consequence of the unavoidable filtering process, the measured phase fluctuation correlations $\rho_{m,n}^{(f)}(x)$ are somewhat lower than the actual ones (see Fig. 16).
@@
Figure 17: Effect of removing phase slips on correlations. The data is for five modes at $350$ mA for lateral position $x=15~{}\mu$m. These are the quintuplet modes shown in Fig. 3. a, Correlation diagram for the phase fluctuations $\delta_{\tau}\phi_{mr}^{(f)}$ of modes $2$ and $4$ before removal of the phase slips and b after their removal. The phase fluctuations are normalized by their standard deviation. The correlation coefficients are $0.05$ and $0.33$, respectively. c, Correlation matrix of the five modes for $x=15~{}\mu$m before removal of phase slips and d after their removal. Please note the different color scales.
@@
Figure 17 shows how strongly the phase slips that occur as artifacts at local minima of the amplitude affect the measured PFCs. The correlation diagram of two modes of a quintuplet at $350$ mA in Fig. 17a has a cross shape which is due to the very strong phase fluctuations that appear during the phase slips. Consequently, the correlation coefficient is only $0.05$. By removing $4$ ns long windows around the phase slips as described above, these spurious fluctuations are removed, and the correlation diagram becomes an elliptic point cloud with a correlation coefficient of $0.33$ as shown in Fig. 17b. The correlation matrices of the quintuplet modes before and after removal of the phase slips are shown in Figs. 17c and d, respectively. The correlations of almost all mode pairs increase drastically when removing the phase slips, and the correlation data shown in Fig. 4 calculated after their removal. Please note that the correlation values in Fig. 17d are different from those in Fig. 4a because the latter includes a spatial average of the PFCs as explained in the following.
@@
Moreover, detector noise and signals from the dynamics of the BAL that fall within the filter bandwidth significantly reduce the measured PFCs as well. The signal-to-noise ratio of the beat note for mode $m$ at position $x$ is given by
@@
where the RF-spectrum $P_{\mathrm{bal}}$ of the BAL is calculated from its time trace $U_{\mathrm{bal}}(t,x)$ in the same way as $P_{hd}$ in Eq. (35). The joint SNR of the mode pair $(m,n)$ is defined as
$\mathrm{SNR}_{m,n,x}=(\mathrm{SNR}_{m,x}+\mathrm{SNR}_{n,x})/2\,,$ (39)
and the abscissa of Figs. 4b and c is simply $\max_{x}(\mathrm{SNR}_{m,n,x})$.
Figure 18a and b show the PFC and the signal-to-noise ratio (SNR) of two modes of 1st and 2nd order as function of the lateral position $x$. Not surprisingly the SNR and in consequence the PFC is low at positions where the modes have low intensities. Figures 18c and d show the correlation diagrams of the PFC and the SNR of the individual modes, which demonstrate how strongly the measured PFCs depend on a good SNR. Hence, we only consider the data points with the best joint SNR, that is,
$\mathrm{SNR}_{m,n,x}\geq\mathrm{SNR}_{th}=\max_{x}(\mathrm{SNR}_{m,n,x})-5\,\mathrm{dB}\,.$ (40)
We calculate $\langle\rho_{m,n}^{(f)}\rangle_{x}$ as the average of $\rho_{m,n}^{(f)}(x)$ over the data points fulfilling this criterion. The SNR threshold is indicated by the cyan line in Fig. 18b, and the corresponding data points are indicated by cyan dots in Fig. 18a. The average of these data points is $\langle\rho_{2,4}^{(f)}\rangle_{x}=0.24\pm 0.05$ (green line in Fig. 18a).
The example in Fig. 18 is typical for mode pairs with fairly strong correlation and decent SNR. Figure 19 shows the example of a NNN-mode pair at $300$ mA which has the highest correlation that was found in the measurement data, $\langle\rho_{4,6}^{(f)}\rangle_{x}=0.76\pm 0.03$. The sum of the linewidths of these lasing modes is $40.7$ MHz compared to a linewidth of $9.2$ MHz of the beat note, which yields a linewidth reduction factor of $F_{\mathrm{red}}=4.4$. For comparison, the measured PFC of $0.76$ corresponds to $F_{\mathrm{red}}=4.14$ according to Eq. (20), which agrees almost quantitatively. We see in Figs. 19c and d that the PFC saturates with increasing SNR, and it seems that this saturation indicates when the correct value of the PFC is reached. However, it should be noted that this example has an unusually low LRF; in most cases the $F_{\mathrm{red}}\geq 10$ (see Fig. 3e), but measuring PFCs above $0.9$ appears to be out of reach with the current setup.

### v2

@@ D.2 Simulations
In this section we present simulations of a heterodyning signal in order to demonstrate how the phase fluctuations can be extracted from it and to understand the effect of applying a frequency filter. We numerically calculate a heterodyning beat note $P_{mr}(t)$ as in Eq. (22) with a sampling step of $\tau=20$ ps where we choose constant mode amplitudes $A_{m}=A_{r}=1$, a beat note frequency $\Delta\nu_{mr}=2$ GHz, and linewidths of $\Gamma_{m}=100$ MHz and $\Gamma_{r}=1$ MHz for the lasing mode and the reference laser, respectively. The phases $\phi_{m}$ and $\phi_{r}$ of lasing mode and reference laser are calculated as in Eq. (8). The beat note $P_{mr}(t)$ is a sinusoidal signal with phase modulated by the phase fluctuations of the two lasing modes.
@@
Figure 18: Analysis of simulated beat note signal. a, Amplitude of the analytic signal of the simulated beat note without filter ($A_{mr}$, blue) and with filter ($A_{mr}^{(f)}$, red). b, Phase extracted from the simulated beat note without filter ($\phi_{mr}$, blue) and with filter ($\phi_{mr}^{(f)}$, red) as well as the original phase ($\phi_{m}-\phi_{r}$, green). c, Original phase fluctuations [$\delta_{\tau}(\phi_{m}-\phi_{r})$, green] and phase fluctuations extracted with filter ($\delta_{\tau}\phi_{mr}^{(f)}$, red).
@@
First, we apply the Hilbert transform directly to $P_{mr}(t)$ and extract the amplitude $A_{mr}(t)$ and the phase $\phi_{mr}(t)$ of the beat note as in Eqs. (24)–(26). The amplitude shown as blue line in Fig. 18 exhibits small fluctuations around $2$ even though it is expected to be constant and equal to $2$. Figure 18b compares the phase $\phi_{mr}(t)$ (blue) extracted from the analytic signal according to Eq. (26) to the original phase $\phi_{m}-\phi_{r}$ (green) used to calculate the beat note, which agree well. This demonstrates that the procedure outlined above allows us to determine the phase fluctuations of the lasing modes from the beat notes, even though there are small deviations between extracted amplitude and phase compared to the simulated ones.
@@
However, in practice the measured heterodyne signals contain several beat notes as well as contributions from the BAL dynamics, hence a relatively narrow filter must be applied around the individual beat notes in order to separate them from the other signals. To test the effect of filtering, we apply a $300$ MHz wide filter [cf. Eq. (34)] to the simulated beat note $P_{mr}(t)$. Then we calculate the analytic signal of the filtered time trace $P_{mr}^{(f)}$ and extract amplitude $A_{mr}^{(f)}$ and phase $\phi_{mr}^{(f)}$ as above, which are shown in red in Fig. 18. The amplitude of the filtered signal shows severe fluctuations, including dips going to almost zero. These fluctuations are artifacts induced by the filter: in the event of very fast phase changes, for example at $46$ ns in Fig. 18b, the instantaneous frequency of the beat note can leave the frequency range of the filter, thus reducing the amplitude of the filtered signal. This unfortunately means that our heterodyne analysis does not allow us to determine a possible fast amplitude dynamics of the lasing modes since any genuine fluctuations of the amplitude will be obscured by the spurious amplitude fluctuations due to the filter.
@@
The phase (Fig. 18b) is also affected by the filter. First, it is smoothed because very fast phase fluctuations are filtered out. Hence the phase fluctuations of the filtered signal $\delta_{\tau}\phi_{mr}^{(f)}$ (red line in Fig. 18c) have a smaller variance than the original ones (green line) and are also smooth, not randomly fluctuating like the original ones. Second, strong dips in the amplitude lead to spurious fluctuations of the phase (phase slips) like at $46$ ns where $\phi_{mr}^{(f)}$ no longer follows the original phase. These phase slips manifest in strong extrema of the phase fluctuations as seen in Fig. 18c at $46$ ns. Once the amplitude has recovered (from $55$ ns on), the filtered phase once again follows the original one, but with an offset. In order to remove these spurious phase fluctuations induced by the filter, $4$ ns long windows around local minima of the amplitude are removed from the phase fluctuation time traces before calculating their correlation (see Appendix D.3).
@@
Finally, we investigate how the data analysis procedure affects the correlation measurement. To this end, we calculate the beat note $P_{nr}$ of another lasing mode $n$ with the reference laser. Lasing modes $n$ and $m$ have the same linewidth of $100$ MHz, but we now introduce a correlation $\rho_{0}$ of their phase fluctuations. The phase fluctuations of mode $n$ are calculated from the random variable $z_{n}$ [see Eq. (8)] given by
@@
where $z_{m}$ and $z_{3}$ are independent Gaussian random variables with unit variance, and $z_{m}$ is the random variable determining the phase fluctuations of mode $m$. By construction, $\rho_{z_{m},z_{n}}=\rho_{0}$.
We calculate the Pearson correlation coefficient $\rho_{m,n}^{(f)}$ of the filtered phase fluctuations $\delta_{\tau}\phi_{mr}^{(f)}$ and $\delta_{\tau}\phi_{nr}^{(f)}$, after additionally removing the parts near local minima of the amplitude as described above. The resulting correlations $\rho_{m,n}^{(f)}$ are compared to the actual correlations $\rho_{0}$ of the simulated phase fluctuations in Fig. 19. We find that for small $\rho_{0}$, the extracted PFCs are $\rho_{m,n}^{(f)}\simeq\rho_{0}/2$ and remain significantly lower even for high $\rho_{0}$. In conclusion, the analysis procedure used to measure the PFCs yields values that are systematically lower than the actual correlations, which partially explains the relatively low PFCs observed experimentally. Nonetheless, our simulations demonstrate that we obtain qualitatively correct information about possible phase correlations between different lasing modes.
@@ D.3 Analysis of experimental data
It should be noted that some lasing modes are ”degenerate” in the RF-spectrum, that is, they have almost the same beat note frequency $f_{m}$ because their optical frequencies are symmetrical with respect to the reference laser frequency. Hence, it is impossible to separate them with a filter as explained in the previous section, and they are left out of the analysis since they would give spurious results.
@@
where $\Gamma_{m}$ is the full width at half maximum (FWHM) of the beat note averaged over $x$, $\Gamma_{m}=\langle\Gamma_{m,x}\rangle_{x}$. Here, the FWHM $\Gamma_{m,x}$ at position $x$ is obtained from the RF-spectrum of the heterodyne signal (smoothed with a $f_{\mathrm{avg}}=10$ MHz moving average),
$P_{hd}(f,x)=\mathcal{M}\{|\tilde{U}(f,x)|^{2}\}\,,$ (36)
by determining the $-3$ dB points around the peak at $f_{m}$. Positions for which the FWHM could not be properly determined due to, e.g., low signal strength, are ignored when calculating $\Gamma_{m}$. A total width of $3\Gamma_{m}$ for the filter is used as a compromise between the requirements of having a large bandwidth and removing noise and interfering signals as much as possible. For a Lorentzian line shape, the filter thus contains $79.5\%$ of the total RF-power of the beat note signal. In practice, the filter width is of the order of a few $100$ MHz.
@@
where $\mathrm{sgn}$ is the sign function. The phase $\phi_{mr}^{(f)}$ is the difference between the instantaneous phases of the lasing mode $m$ and the reference laser. In the following we consider the phase fluctuations
@@
The phase fluctuations $\delta_{\tau}\phi_{mr}^{(f)}(t,x)$ are the difference between the phase fluctuations of the lasing mode and that of the reference laser, where the latter are negligible compared to the former since the linewidth of the reference laser is very small compared to that of the lasing modes in our case (see Appendix D.1). It should be noted that the standard deviation of $\delta_{\tau}\phi_{mr}^{(f)}(t,x)$ is significantly reduced compared to that of the actual phase fluctuations due to the filter that is applied (see Fig. 18).
@@
Furthermore, very fast phase fluctuations lead to high instantaneous frequencies that cause the signal to (partially) leave the frequency domain of the boxcar filter that is applied (see Fig. 18). In such cases, $\delta_{\tau}\phi_{mr}^{(f)}$ shows spurious values, and the instantaneous amplitude $A_{mr}^{(f)}$ exhibits a strong dip. To remove these filter artifacts, $4$ ns long pieces of $\delta_{\tau}\phi_{mr}^{(f)}$ are removed around pronounced dips in the amplitude. Finally, the Pearson correlation coefficient $\rho_{m,n}^{(f)}(x)$ between these reduced versions of $\delta_{\tau}\phi_{m}^{(f)}(t,x)$ and $\delta_{\tau}\phi_{n}^{(f)}(t,x)$ is calculated. As a consequence of the unavoidable filtering process, the measured phase fluctuation correlations $\rho_{m,n}^{(f)}(x)$ are somewhat lower than the actual ones (see Fig. 19).
@@
Figure 20: Effect of removing phase slips on correlations. The data is for five modes at $350$ mA for lateral position $x=15~{}\mu$m. These are the quintuplet modes shown in Fig. 3. a, Correlation diagram for the phase fluctuations $\delta_{\tau}\phi_{mr}^{(f)}$ of modes $2$ and $4$ before removal of the phase slips and b after their removal. The phase fluctuations are normalized by their standard deviation. The correlation coefficients are $0.05$ and $0.33$, respectively. c, Correlation matrix of the five modes for $x=15~{}\mu$m before removal of phase slips and d, after their removal. Please note the different color scales.
@@
Figure 20 shows how strongly the phase slips that occur as artifacts at local minima of the amplitude affect the measured PFCs. The correlation diagram of two modes of a quintuplet at $350$ mA in Fig. 20a has a cross shape which is due to the very strong phase fluctuations that appear during the phase slips. Consequently, the correlation coefficient is only $0.05$. By removing $4$ ns long windows around the phase slips as described above, these spurious fluctuations are removed, and the correlation diagram becomes an elliptic point cloud with a correlation coefficient of $0.33$ as shown in Fig. 20b. The correlation matrices of the quintuplet modes before and after removal of the phase slips are shown in Figs. 20c and d, respectively. The correlations of almost all mode pairs increase drastically when removing the phase slips, and all correlation data shown in the main text is calculated after their removal. Please note that the correlation values in Fig. 20d are different from those in Fig. 4a because the latter includes a spatial average of the PFCs as explained in the following.
@@
Moreover, detector noise and signals from the dynamics of the BAL that fall within the boxcar filter bandwidth significantly reduce the measured PFCs as well (see Fig. 21). The signal-to-noise ratio of the beat note for mode $m$ at position $x$ is given by
@@
where the RF-spectrum $P_{\mathrm{BAL}}$ of the BAL alone is calculated from its time trace $U_{\mathrm{BAL}}(t,x)$ in the same way as $P_{hd}$. The joint SNR of the mode pair $(m,n)$ is defined as their geometric mean (or arithmetic mean in logarithmic scale),
$\mathrm{SNR}_{m,n,x}=\sqrt{\mathrm{SNR}_{m,x}\cdot\mathrm{SNR}_{n,x}}\,,$ (40)
and the abscissa of Figs. 4b and c is simply $\max\limits_{x}(\mathrm{SNR}_{m,n,x})$.
Figures 21a and b show the PFC and the signal-to-noise ratio (SNR) of two modes with $q=1$ and $2$ as function of the lateral position $x$. Not surprisingly, the SNR and in consequence the PFC is low at positions where the modes have low intensities. Figures 21c and d show the correlation diagrams of the PFC and the SNR of the individual modes, which demonstrate how strongly the measured PFCs depend on a good SNR. Hence, we only consider the data points with the best joint SNR, that is,
$\mathrm{SNR}_{m,n,x}>\mathrm{SNR}_{th}=\max_{x}(\mathrm{SNR}_{m,n,x})-5\,\mathrm{dB}\,.$ (41)
We calculate $\langle\rho_{m,n}^{(f)}\rangle_{x}$ as the average of $\rho_{m,n}^{(f)}(x)$ over the data points fulfilling this criterion. It is the quantity $\langle\rho_{m,n}^{(f)}\rangle_{x}$ which is shown in Fig. 4. The SNR threshold is indicated by the cyan line in Fig. 21b, and the corresponding data points are indicated by cyan dots in Fig. 21a. The average of these data points is $\langle\rho_{2,4}^{(f)}\rangle_{x}=0.24\pm 0.05$ (green line in Fig. 21a).
The example in Fig. 21 is typical for mode pairs with fairly strong correlation and decent SNR. Figure 22 shows the example of a NNN-mode pair at $300$ mA which has the highest correlation that was found in the measurement data, $\langle\rho_{4,6}^{(f)}\rangle_{x}=0.76\pm 0.03$. The sum of the linewidths of these lasing modes is $40.7$ MHz compared to a linewidth of $9.2$ MHz of the beat note, which yields a linewidth reduction factor of $F_{\mathrm{red}}=4.4$. For comparison, the measured PFC of $0.76$ corresponds to $F_{\mathrm{red}}=4.14$ according to Eq. (21), which agrees almost quantitatively. We see in Figs. 22c and d that the PFC saturates with increasing SNR, and it seems that this saturation indicates when the correct value of the PFC is reached. However, it should be noted that this example has an unusually low LRF; in most cases the $F_{\mathrm{red}}\geq 10$ (see Fig. 3e), but measuring PFCs above $0.9$ appears to be out of reach with the current setup due to insufficient SNR.